OpenStudy (anonymous):

the sum of three numbers is 55. the second number is eight less than the first number and the third number is three more than twice the first number. what are the numbers?

OpenStudy (gorv):

let first number= x

OpenStudy (gorv):

second=x-8

OpenStudy (gorv):

asit is 8 less thn first

OpenStudy (gorv):

3rd number =2*x+3 3 more than twice of first

OpenStudy (gorv):

sum=x+x-8+2x+3=55

OpenStudy (gorv):

sum it up and do further

OpenStudy (gorv):

welcome to OS giel :)

OpenStudy (anonymous):

thankyou :)